OpenStudy (anonymous):

Find the function rule! I'll give a medal/fan!!!! x: -2, -1, 0, 1, 2 y: 9, 4, -1, -6, -11

OpenStudy (anonymous):

Take your time.

OpenStudy (anonymous):

Paper an pencil helps

OpenStudy (anonymous):

Is the slope constant?

OpenStudy (anonymous):

Don't know, all it says is "Find the function" And then it gives me the numbers... That's it

OpenStudy (anonymous):

It's a line dude.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

When x goes up by one, how much does y change by?

OpenStudy (anonymous):

It goes down by 5

OpenStudy (anonymous):

slope is therefore -5

OpenStudy (anonymous):

use \(y=mx+b\)

OpenStudy (anonymous):

Now we know m=-5

OpenStudy (anonymous):

So the function is???

OpenStudy (anonymous):

?

OpenStudy (anonymous):

All I know is x goes up by 1 every time and Y goes down 5.

OpenStudy (anonymous):

Help??????????????????????????????????????????????!!!!

OpenStudy (anonymous):

Ummm What else do you know?

OpenStudy (anonymous):

do you know the equation for a line?

OpenStudy (anonymous):

Not that far in math..

OpenStudy (anonymous):

Ok, here is the rule then: \[ f(x+1) = f(x) - 5 \]

OpenStudy (anonymous):

I got to go :/ Ill give you the fan/ medal anyways Thanks

OpenStudy (anonymous):

Then we say: \[ f(x+n) = f(x) -\underbrace{ 5 - 5 - \ldots -5}_n = f(x) - 5n \]

OpenStudy (anonymous):

If we let \(x=0\), then: \[ f(n) = f(0) - 5n \]

OpenStudy (anonymous):

Also, when \(x=0\), then \(y = f(x) = f(0) =-1\).

OpenStudy (anonymous):

So \[ f(n) = -1-5n \]